Calculate each of the following quantities in 2.50 moles H3PO4:

atoms of O
1
Expert's answer
2020-10-14T05:55:26-0400


Given,

the molicular formula = H3PO4


Hence, no. Of 'O' atom in each molicule of H3PO4 is = 4




1 mole = 6.023×10^23 molicules


2.50 mole =( 2.50× 6.023×10^23 )molicules


= 15.057 × 10^23 molicules




So,


In 1 molicule of H3PO4, 4 atom of O


In (15.057 × 10^23) molicules of H3PO4, ( 4×15.057 × 10^23) atoms of O



Hence,

in 2.50 moles of H3PO4 there are 60.228 × 10^23 atoms of O